OpenStudy (anonymous):

Solve each of the quadratic equations below. @amistre64 x2 − 16 = 0 and x2 = −2x + 24

OpenStudy (nurali):

x^2 - 16 = 0 (x-4)(x+4) = 0 x = 4 or x = -4

OpenStudy (harsimran_hs4):

x^2 = 16 x = +4, -4 equation 2 x^2 + 2x -24 = 0 can you solve this equation ?

OpenStudy (nurali):

x^2 + 2x - 24 = 0 x^2+6x+4x-24=0 (x + 6)(x - 4) = 0 (x + 6)=0 or (x - 4)=0 x =-6 or x = 4